There is a cave accessible only by water. The wate
Answer #1

density of air=1.225kg/m3

pressure=density*force of gravity*height of the air

height of the air=15-10=5m

pressure=1.225*9.8*5=60.025 pa

pressure=60.025/24=2.5 atm